Evaluate

x^2y^3 + x(

1 + y^2)dy/dx= 0




1
Expert's answer
2021-07-20T10:06:33-0400
"x^2y^3+x(1+y^2)\dfrac{dy}{dx}=0"

"x=0, y\in\R"


"y=0, x\in \R"


"xy\not=0"

"\dfrac{1+y^2}{y^3}dy=-xdx"

"\int \dfrac{1+y^2}{y^3}dy=-\int xdx"

"-\dfrac{1}{y^2}+\ln|y|=-\dfrac{1}{2}x^2+C"

"-\dfrac{1}{y^2}+\ln|y|+\dfrac{1}{2}x^2=C"